You place $2000 in an aggressive bank account. How much will be in your account after 21 years if your money grows by 45% every 3 years?
Answer by lwsshak3(6505)

You place $2000 in an aggressive bank account. How much will be in your account after 21 years if your money grows by 45% every 3 years
..
compound interest formula: A=P(1+r)^n, P=initial investment, r=interest rate per period. n=number of periods, A= initial investment after n years
..
for given problem:
Initial investment=2000
nterest rate per period=45%
number of periods=21/3=7
A=2000(1+.45)^7
=2000(1.45)^7=26953
ans:
After 21 years, $26,953, will be in your account